OpenStudy (anonymous):

Solve x^2-3x-10=0 by completing the square.

OpenStudy (anonymous):

(x-5)(x+2) = 0 x = 5, -2

OpenStudy (anonymous):

The answer is correct. Is that good enough for you? Or do you want to see how it was done?

OpenStudy (anonymous):

Can I see how it was done?

OpenStudy (anonymous):

i pretty much did all the work You could add this step: x^2 - 5x + 2x - 10 = 0

OpenStudy (anonymous):

and then the rest

OpenStudy (anonymous):

OK, give me a moment to write all the steps out.

OpenStudy (anonymous):

\[Ax ^{2}+Bx+C=0\] First step take the B value and divide it by 2. So you have 3/2. Second step square the value you found in step 1. So you have 9/4 Third step rewrite the equation. So you have \[x ^{2}-3x-9/4+9/4-10=0\]This will factor out to be \[(x-3/2)^{2}-9/4-10=0\] This becomes \[(x-3/2)^{2}-49/4\] This becomes\[(x-3/2)^{2}=49/4\]Now that the square root of both sides.\[(x-3/2)=\pm7/2\]\[x=\pm7/2+3/2\] x=-2 or x=5

OpenStudy (anonymous):

Thank you! :D

OpenStudy (anonymous):

When I wrote divide the B value is should be -3/2. And it should say Now take the square root of both sides.